# EvoP: Robust LLM Inference via Evolutionary Pruning

### v1

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success in natural language processing tasks, but their massive size and computational demands hinder their deployment in resource-constrained environments. Existing structured pruning methods address this issue by removing redundant structures (e.g., elements, channels, layers) from the model. However, these methods employ a heuristic pruning strategy, which leads to suboptimal performance. Besides, they also ignore the data characteristics when pruning the model.
To overcome these limitations, we propose EvoP, an evolutionary pruning framework for robust LLM inference. EvoP first presents a cluster-based calibration dataset sampling (CCDS) strategy for creating a more diverse calibration dataset. EvoP then introduces an evolutionary pruning pattern searching (EPPS) method to find the optimal pruning pattern. Compared to existing structured pruning techniques, EvoP achieves the best performance while maintaining the best efficiency. Experiments across different LLMs and different downstream tasks validate the effectiveness of the proposed EvoP, making it a practical and scalable solution for deploying LLMs in real-world applications.
EvoP: Robust LLM Inference via Evolutionary Pruning
Shangyu Wu CityU                        Hongchao Du CityU                        Ying Xiong thanks: Corresponding author. MBZUAI                        Shuai Chen Baidu
Tei-Wei Kuo National Taiwan University                        Nan Guan CityU                        Chun Jason Xue MBZUAI
@@ 1 Introduction
Large Language Models (LLMs), such as Llama Touvron et al. (2023b) and OPT Zhang et al. (2022a), have revolutionized natural language processing (NLP) by achieving state-of-the-art performance across a wide range of tasks Touvron et al. (2023a); Zhang et al. (2022a); Scao et al. (2022); Chowdhery et al. (2023); Thoppilan et al. (2022); Zeng et al. (2023); Chowdhery et al. (2023). However, the success of these models comes at a significant computational cost. LLM consists of billions of parameters, requiring substantial memory, storage, and energy resources. This computational burden limits their deployment in resource-constrained environments.
To address these challenges, network pruning LeCun et al. (1989) has emerged as a critical technique for optimizing LLMs by reducing their size while preserving their performance. Pruning involves selectively removing redundant or less important parameters (e.g., elements Sun et al. (2024), channels Ashkboos et al. (2024), or layers Song et al. (2024)) from a pre-trained model, which enables faster inference and lower energy consumption. Importantly, pruning is feasible in LLMs because these models are often over-parameterized, containing more parameters than necessary to achieve their performance.
@@
To address the challenges, we follow layer-wise pruning and introduce a novel evolutionary pruning framework, EvoP. EvoP first presents a cluster-based calibration dataset sampling strategy for collecting a more diverse calibration dataset. EvoP then combines the diverse calibration dataset with an evolutionary pruning pattern search method to identify the optimal pruning patterns that generalize well across tasks. Extensive experiments demonstrate that the proposed EvoP outperforms existing state-of-the-art pruning methods of different granularity. The proposed EvoP also performs well on in-domain and out-of-domain perplexity datasets, showing its better generalization capabilities. Codes are available at  [^1] https://anonymous.4open.science/r/EvoP-D4DA.
@@ 2.1 Network Pruning
Network pruning is a key technique in the LLM era, which aims to reduce neural networks’ size and computational complexity by removing unnecessary or redundant structures, such as elements Sun et al. (2024), channels Ashkboos et al. (2024), or layers Song et al. (2024). Existing methods target different levels of pruning granularity, from fine-grained element-wise pruning to coarse-grained layer-wise removal, offering a trade-off between model compression, hardware efficiency, and task performance.
@@
, where $\mathcal{L}(\cdot)$ is the loss function, $p\in\{0,1\}^{m}$ is a binary vector of zeros and ones, indicating a pruning pattern over $m$ pruning components, which can be weight elements, channels, or layers. $p[i]=1$ represents the $i$-th components in the model $\mathcal{M}$ is pruned, and vice versa. Moreover, $f(,)$ is the function that applies the pruning pattern $p$ to the pre-trained model $\mathcal{M}$ with different pruning techniques. $\|p\|_{0}$ and $\|W\|_{0}$ are the number of non-zeros values in $p$ and the total number of parameters in $W$.
Figure 1: Perplexity loss of the Tinyllama model on the calibration dataset across different sparsity. The Ideal is the globally optimal solution in each pruning pattern space. The Gene is the solution using the genetic algorithm. SLEB fails to find optimal when the pruning pattern space becomes larger (large sparsity).
Figure 2: Visualization of the semantic distribution of the calibration dataset. Each point represents an input sentence. Points with the same color are clustered together by KMeans.
@@ 2.2 Challenges of Different Pruning Techniques
The above formulation provides a unified framework for understanding and comparing different pruning methods. However, each pruning granularity introduces unique challenges. Element-wise pruning is a kind of unstructured pruning, which is not a hardware-friendly technique, requiring specific computation kernel designs and implementations. For example, 2:4 pruning Frantar and Alistarh (2023); Sun et al. (2024); Zhang et al. (2024b); Zhou et al. (2021); Zhang et al. (2022b) relies on specific sparse matrix multiplication operators and is currently only supported by GPUs with NVIDIA Ampere architecture. Besides, the pruning pattern space $P$ of the element-wise pruning can be enormous, resulting in the proposed solutions being generally heuristic algorithms on each weight matrix (Wang, 2020; Jiang et al., 2018).
Channel-wise pruning removes the entire channel for better hardware efficiency (Ashkboos et al., 2024; Ma et al., 2023). However, complex dependencies between channels often make it challenging to determine which channel should be pruned, thus leading to substantial performance drops Voita et al. (2019).
Recent works Huang et al. (2024); Song et al. (2024) pay more attention to coarse-grained pruning, i.e., layer-wise pruning, due to its greater hardware efficiency and better model performance. Those works can be categorized into two branches: early exit and layer dropping. The former introduces intermediate exit points within the model and terminates the inference at the predicted exit point (Huang et al., 2024; Xin et al., 2020), while the latter drops some layers based on the calibration dataset and leaves the pruned model for the inference (Song et al., 2024). Compared to layer dropping, although early exit seems an easier-to-solve approach to pruning the model, it isn’t easy to accelerate the inference in batches and reduce memory overheads in deployment Song et al. (2024). Considering the above challenges of different pruning techniques, this paper mainly focuses on optimizing the layer dropping in layer-wise pruning.
@@ 3 Observations
Figure 3: Perplexity of SLEB and Gene using different calibration datasets. The sparsity of Tinyllama is set to 50% for large enough PPS. ‘C0’, ‘C1’, …, and ‘C4’ are randomly sampled within each cluster.
@@ 3.1 Observation 1: Sub-optimal Pruning Pattern of Heuristics in Large PPS
Existing state-of-the-art layer-dropping techniques adopt heuristic algorithms to select layers to drop. For example, SLEB Song et al. (2024) enumerates all unpruned layers and removes the one where the model without the layer achieves the minimum perplexity loss on the calibration dataset. Then, SLEB repeats this process until the given sparsity is reached. This simple yet effective technique is successful because the SLEB algorithms can achieve as optimal as ideal solutions when the pruning pattern space is not large enough (small sparsity).
To better support the above claims, we conducted an analysis experiment on the Tinyllama model Zhang et al. (2024a) over the calibration dataset in Figure 1. We follow the data selection of SLEB to build the calibration dataset from WikiText-2 Merity et al. (2017). We collect three sets of perplexity losses over different sparsities: the SLEB, the Ideal that enumerates the whole pruning pattern space, and the Gene using the genetic algorithm to search.
As expected, when the sparsity is below around 35%, SLEB can easily find the global optimal since the pruning pattern space is small, only seven layers need to be selected at most. However, when the size of the pruning pattern space increases (larger sparsity), SLEB gradually gets stuck in the local optimal. Therefore, this motivates us to propose a new algorithm to further optimize the pruning in the scenarios with larger pruning pattern space. Noted, the size of the pruning pattern space is not only related to the given sparsity but also depends on the original model size, which can be calculated as the combinatorial number  [^2] The number of model layers is the total number of $N$, and the number of pruned layers is the degree $k$, thus the size of the pruning pattern space is $\binom{N}{k}$.. Therefore, even a smaller sparsity (e.g., 10%-20%) can lead to an ample pruning pattern space for larger models.
@@ 3.2 Observation 2: Low Data Diversity of Calibration Dataset
The above analysis shows the potential to approach the global optimal. However, the global optimal might differ when pruning on different calibration datasets. First, we analyze the semantic distribution of WikiText-2 and visualize the distribution in Figure 2. We use the BERT-base Devlin et al. (2019) model to encode each sentence and TSNE to reduce the embedding dimensions. Figure 2 demonstrates that some sentences are semantically similar and naturally clustered. We also use KMeans to cluster those sentences and draw the colors, proving this claim.
Based on the data characteristics, we further show the impact of different calibration datasets on the global optimal pruning patterns. Since Figure 1 has demonstrated the Gene can achieve comparable performance to the Ideal, we would regard the Gene’s results as the Ideal’s. We collect the perplexity of SLEB and Gene using the samples from each cluster in Figure 3. Experimental results show that different calibration datasets may lead to different optimal. This motivates us to propose a new data sampling method to help the pruning algorithms reach a better optimal.
@@ 4.1 CCDS: Cluster-based Calibration Dataset Sampling
To achieve a more robust pruning pattern, EvoP first partition the raw dataset (e.g., WikiText-2 Merity et al. (2017) or C4 Raffel et al. (2020)) into $k$ clusters based on input semantic similarity. To ensure the clustering effect and semantic integrity, we first divide the dataset into chunks with the same number of sentences and then cluster the data at the chunk level. BERT Devlin et al. (2019) is applied to generate the embedding of each chunk since it can get information-rich representations compared to unidirectional pre-trained language models. Then, clustering algorithms such as KMeans are used to cluster these embeddings.
The generated clusters separate inputs with different computational requirements, allowing for more diverse and representative data sampling. For each cluster $C_{j}$ (where j = 1,2$\dots$,k), We randomly select $n$ representative samples, each of which may contain multiple chunks. These samples are used to evaluate the performance of different pruning patterns during the search process. Sampling across all clusters ensures data diversity, improving the robustness of the pruning patterns. The detailed algorithm is shown in Algorithm 4.1.
@@
0:  Raw Dataset $\mathcal{D}$, number of clusters $k$, number of samples per cluster $n$, the maximum length of each samples $len$
@@
5:  $\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{k}\}=\text{KMeans}(E,k)$, where $E=\{e_{1},e_{2},\dots,e_{max}\}$
@@ 4.2 EPPS: Evolutionary Pruning Pattern Search
Unlike heuristic methods in SLEB Song et al. (2024), which rely on predefined rules (e.g., magnitude-based pruning), we proposed a revolutionary pruning pattern search algorithm (Algorithm 4.2) to explore search space of possible pruning patterns systematically and adaptively. First, a population of pruning patterns is initialized, where each pattern is represented as a binary vector $p\in\{0,1\}^{m}$, with $m$ is the number of layers in the LLM. A value of 1 indicates that the layer is pruned, while 0 means it is retained. Then, for each pruning pattern $p$, the fitness is computed as the average loss on the $k\times n$ representative samples. The loss is measured using the task-specific objective (e.g., cross-entropy loss for language modeling). EvoP uses perplexity to measure the loss and select the top patterns with the lowest loss (highest fitness) to generate the next generation [^3] We select top 30% in the implementation.. This mimics natural selection, where the fittest individuals survive. New pruning patterns are generated by combining (crossover) and randomly altering (mutation) the selected patterns. This introduces diversity into the population, enabling the exploration of new pruning configurations. Note that the newly pruning patterns may not meet the sparsity requirements, so we randomly select several corresponding positions to flip until the required sparsity is met (double mutation). Last, the process repeats for several generations or until convergence. The final pruning pattern for the cluster is the one with the lowest loss.
@@ 5 Experiments
5.1 Datasets and Experiments Setups
Datasets. Following SLEB Song et al. (2024), we also use the library LM-Harness Gao et al. (2024) to evaluate our EvoP and other pruning techniques on the same five representative datasets. These datasets cover a wide range of downstream tasks, including reasoning, commonsense understanding, and question answering, making them suitable for evaluating the generalization ability of pruned models.
Models. We conducted experiments on three large language models, i.e., Llama-2-13b Touvron et al. (2023b), OPT-13b, and OPT-30b Zhang et al. (2022a). These models can serve as representative benchmarks to analyze the effectiveness of our EvoP across different architectures and scales.
@@ 5.1 Datasets and Experiments Setups
  • Wanda Sun et al. (2024): A superior element-wise pruning that prunes weights with the smallest magnitudes multiplied by the corresponding input activations.
  • SliceGPT Ashkboos et al. (2024): An advanced channel-wise pruning that prunes the columns or rows of weights.
  • SLEB Song et al. (2024): A surpassing layer-wise pruning that removes some layers or blocks of the model.
@@
Implementation Details. The proposed EvoP was implemented using the PyTorch and Transformer frameworks. We evaluated the baseline with Llama-2-13b and OPT-13b on one NVIDIA A100 GPU with 40GB memory and the baseline with OPT-30b on one NVIDIA A800 GPU with 80GB memory. The size of the calibration dataset includes 5 samples with 2048 tokens each, which is equivalent to the dataset used in SLEB Song et al. (2024). The number of clusters in EvoP is set to 5, and the max number of iterations of EPPS is set to 100 for models with 13 billion parameters and 200 for models with 30 billion parameters.

### v2

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success in natural language processing tasks, but their massive size and computational demands hinder their deployment in resource-constrained environments. Existing model pruning methods address this issue by removing redundant structures (e.g., elements, channels, layers) from the model. However, these methods employ a heuristic pruning strategy, which leads to suboptimal performance. Besides, they also ignore the data characteristics when pruning the model.
To overcome these limitations, we propose EvoP, an evolutionary pruning framework for robust LLM inference. EvoP first presents a cluster-based calibration dataset sampling (CCDS) strategy for creating a more diverse calibration dataset. EvoP then introduces an evolutionary pruning pattern searching (EPPS) method to find the optimal pruning pattern. Compared to existing model pruning techniques, EvoP achieves the best performance while maintaining the best efficiency. Experiments across different LLMs and different downstream tasks validate the effectiveness of the proposed EvoP, making it a practical and scalable solution for deploying LLMs in real-world applications.
Keywords:
Large Language Model Layer-wise Model Pruning Evolutionary Search.
@@ 1 Introduction
Large Language Models (LLMs), such as Llama-series [16] and OPT-series [23], have revolutionized natural language processing (NLP) by achieving state-of-the-art performance across a wide range of tasks [17, 23, 12, 2, 15, 21]. However, the success of these models comes at a significant computational cost. LLM consists of billions of parameters, requiring substantial memory, storage, and energy resources. This computational burden limits their deployment in resource-constrained environments.
To address these challenges, network pruning [8] has emerged as a critical technique for optimizing LLMs by reducing their size while preserving their performance. Pruning involves selectively removing redundant or less important parameters (e.g., elements [14], channels [1], or layers [13]) from a pre-trained model, which enables faster inference and lower energy consumption. Importantly, pruning is feasible in LLMs because these models are often over-parameterized, containing more parameters than necessary to achieve their performance.
@@
To address the challenges, we follow layer-wise pruning and introduce a novel evolutionary pruning framework, EvoP. EvoP first presents a cluster-based calibration dataset sampling strategy for collecting a more diverse calibration dataset. EvoP then combines the diverse calibration dataset with an evolutionary pruning pattern search method to identify the optimal pruning patterns that generalize well across tasks. Extensive experiments demonstrate that the proposed EvoP outperforms existing state-of-the-art pruning methods of different granularity. The proposed EvoP also performs well on in-domain and out-of-domain perplexity datasets, showing its better generalization capabilities. Codes are available at https://github.com/luffy06/EvoP.
@@ 2.1 Network Pruning
Network pruning is a key technique in the LLM era, which aims to reduce neural networks’ size and computational complexity by removing unnecessary or redundant structures, such as elements [14], channels [1], or layers [13]. Existing methods target different levels of pruning granularity, from fine-grained element-wise pruning to coarse-grained layer-wise removal, offering a trade-off between model compression, hardware efficiency, and task performance.
@@
where $\mathcal{L}(\cdot)$ is the loss function, $p\in\{0,1\}^{m}$ is a binary vector of zeros and ones, indicating a pruning pattern over $m$ pruning components, which can be weight elements, channels, or layers. $p[i]=1$ represents the $i$-th components in the model $\mathcal{M}$ is pruned, and vice versa. Moreover, $f(,)$ is the function that applies the pruning pattern $p$ to the pre-trained model $\mathcal{M}$ with different pruning techniques. $\|p\|_{0}$ and $\|W\|_{0}$ are the number of non-zeros values in $p$ and the total number of parameters in $W$.
Figure 1: Observations. (a) Perplexity loss of the Tinyllama model on the calibration dataset across different sparsity. The Ideal is the globally optimal solution in each pruning pattern space. The Gene is the solution using the genetic algorithm. SLEB fails to find optimal when the pruning pattern space becomes larger (large sparsity). (b) Visualization of the semantic distribution of the calibration dataset. Each point represents an input sentence. Points with the same color are clustered together by KMeans. (c) Perplexity of SLEB and Gene using different calibration datasets. The sparsity of Tinyllama is set to 50% for large enough PPS. ‘C0’, ‘C1’, …, and ‘C4’ are randomly sampled within each cluster.
@@ 2.2 Challenges of Different Pruning Techniques
The above formulation provides a unified framework for understanding and comparing different pruning methods. However, each pruning granularity introduces unique challenges. Element-wise pruning is a kind of unstructured pruning, which is not a hardware-friendly technique, requiring specific computation kernel designs and implementations. For example, 2:4 pruning [4, 14, 25, 26, 24] relies on specific sparse matrix multiplication operators and is currently only supported by GPUs with specific architectures, such as NVIDIA Ampere architecture, Hopper. Besides, the pruning pattern space $P$ of the element-wise pruning can be enormous, resulting in the proposed solutions being generally heuristic algorithms on each weight matrix [19, 7].
Channel-wise pruning removes the entire channel for better hardware efficiency [1, 9]. However, complex dependencies between channels often make it challenging to determine which channel should be pruned, thus leading to substantial performance drops [18].
Recent works [6, 13] pay more attention to coarse-grained pruning, i.e., layer-wise pruning, due to its greater hardware efficiency and better model performance. Those works can be categorized into two branches: early exit and layer dropping. The former introduces intermediate exit points within the model and terminates the inference at the predicted exit point [6, 20], while the latter drops some layers based on the calibration dataset and leaves the pruned model for the inference [13]. Compared to layer dropping, although early exit seems an easier-to-solve approach to pruning the model, it isn’t easy to accelerate the inference in batches and reduce memory overheads in deployment [13]. Considering the above challenges of different pruning techniques, this paper mainly focuses on optimizing the layer dropping in layer-wise pruning.
@@ 3.1 Observation 1: Sub-optimal Pruning Pattern of Heuristics in Large PPS
Existing state-of-the-art layer-dropping techniques adopt heuristic algorithms to select layers to drop. For example, SLEB [13] enumerates all unpruned layers and removes the one where the model without the layer achieves the minimum perplexity loss on the calibration dataset. Then, SLEB repeats this process until the given sparsity is reached. This simple yet effective technique is successful because the SLEB algorithms can achieve as optimal as ideal solutions when the pruning pattern space is not large enough (small sparsity).
To better support the above claims, we conducted an analysis experiment on the Tinyllama model [22] over the calibration dataset in Figure 1(a). We follow the data selection of SLEB to build the calibration dataset from WikiText-2 [10]. We collect three sets of perplexity losses over different sparsities: the SLEB, the Ideal that enumerates the whole pruning pattern space, and the Gene using the genetic algorithm to search.
As expected, when the sparsity is below around 35%, SLEB can easily find the global optimal since the pruning pattern space is small, only seven layers need to be selected at most. However, when the size of the pruning pattern space increases (larger sparsity), SLEB gradually gets stuck in the local optimal. Therefore, this motivates us to propose a new algorithm to further optimize the pruning in the scenarios with larger pruning pattern space. Noted, the size of the pruning pattern space is not only related to the given sparsity but also depends on the original model size, which can be calculated as the combinatorial number. The number of model layers is the total number of $N$, and the number of pruned layers is the degree $k$, thus the size of the pruning pattern space is $\binom{N}{k}$. Therefore, even a smaller sparsity (e.g., 10%-20%) can lead to an ample pruning pattern space for larger models.
@@ 3.2 Observation 2: Low Data Diversity of Calibration Dataset
The above analysis shows the potential to approach the global optimal. However, the global optimal might differ when pruning on different calibration datasets. First, we analyze the semantic distribution of WikiText-2 and visualize the distribution in Figure 1(b). We use the BERT-base [3] model to encode each sentence and TSNE to reduce the embedding dimensions. Figure 1(b) demonstrates that some sentences are semantically similar and naturally clustered. We also use KMeans to cluster those sentences and draw the colors, proving this claim.
Based on the data characteristics, we further show the impact of different calibration datasets on the global optimal pruning patterns. Since Figure 1(a) has demonstrated the Gene can achieve comparable performance to the Ideal, we would regard the Gene’s results as the Ideal’s. We collect the perplexity of SLEB and Gene using the samples from each cluster in Figure 1(c). Experimental results show that different calibration datasets may lead to different optimal. This motivates us to propose a new data sampling method to help the pruning algorithms reach a better optimal.
@@ 4.1 CCDS: Cluster-based Calibration Dataset Sampling
To achieve a more robust pruning pattern, EvoP first partition the raw dataset (e.g., WikiText-2 [10] or C4 [11]) into $k$ clusters based on input semantic similarity. To ensure the clustering effect and semantic integrity, we first divide the dataset into chunks with the same number of sentences and then cluster the data at the chunk level. BERT [3] is applied to generate the embedding of each chunk since it can get information-rich representations compared to unidirectional pre-trained language models. Then, clustering algorithms such as KMeans are used to cluster these embeddings.
[…]
@@
0: Raw Dataset $\mathcal{D}$, number of clusters $k$, number of samples per cluster $n$, the maximum length of each samples lenlen
@@
[…]
@@ 4.2 EPPS: Evolutionary Pruning Pattern Search
Unlike heuristic methods in SLEB [13], which rely on predefined rules (e.g., magnitude-based pruning), we proposed a revolutionary pruning pattern search algorithm (Algorithm 4.2) to explore search space of possible pruning patterns systematically and adaptively. First, a population of pruning patterns is initialized, where each pattern is represented as a binary vector $p\in\{0,1\}^{m}$, with $m$ is the number of layers in the LLM. A value of 1 indicates that the layer is pruned, while 0 means it is retained. Then, for each pruning pattern $p$, the fitness is computed as the average loss on the $k\times n$ representative samples. The loss is measured using the task-specific objective (e.g., perplexity loss for language modeling). EvoP uses perplexity to measure the loss and select the top patterns with the lowest loss (highest fitness) to generate the next generation. We select the top 30% in the implementation. This mimics natural selection, where the fittest individuals survive. New pruning patterns are generated by combining (crossover) and randomly altering (mutation) the selected patterns. This introduces diversity into the population, enabling the exploration of new pruning configurations. Note that the newly pruning patterns may not meet the sparsity requirements, so we randomly select several corresponding positions to flip until the required sparsity is met (double mutation). Last, the process repeats for several generations or until convergence. The final pruning pattern for the cluster is the one with the lowest loss.
@@ 5 Experiments
5.1 Datasets and Experimental Setup
Datasets. Following SLEB [13], we also use the library LM-Harness [5] to evaluate our EvoP and other pruning techniques on the same five representative datasets. These datasets cover a wide range of downstream tasks, including reasoning, commonsense understanding, and question answering, making them suitable for evaluating the generalization ability of pruned models.
Models. We conducted experiments on three large language models, i.e., Llama-2-13b [16], OPT-13b, and OPT-30b [23]. These models can serve as representative benchmarks to analyze the effectiveness of our EvoP across different architectures and scales.
@@ 5.1 Datasets and Experimental Setup
  • Wanda [14]: A superior element-wise pruning that prunes weights with the smallest magnitudes multiplied by the corresponding input activations.
  • SliceGPT [1]: An advanced channel-wise pruning that prunes the columns or rows of weights.
  • SLEB [13]: A surpassing layer-wise pruning that removes some layers or blocks of the model.
@@
Implementation Details. The proposed EvoP was implemented using the PyTorch and Transformer frameworks. We evaluated the baseline with Llama-2-13b and OPT-13b on one NVIDIA A100 GPU with 40GB memory and the baseline with OPT-30b on one NVIDIA A800 GPU with 80GB memory. The size of the calibration dataset includes 5 samples with 2048 tokens each, which is equivalent to the dataset used in SLEB [13]. The number of clusters in EvoP is set to 5, and the max number of iterations of EPPS is set to 100 for models with 13 billion parameters and 200 for models with 30 billion parameters.